OpenStudy (anonymous):

Need help rewriting the equation so y is a function of x in the problem 3x+2y=-7

OpenStudy (anonymous):

Well I just write it in slope intercept form first. 3x+2y=-7 3x-3x+2y=-7-3x 2y=-3x-7 /2 /2 /2 y = -3/2x - 7/2

OpenStudy (anonymous):

Thank you so much. That really helped out.